OpenStudy (anonymous):

1.) y = –x + 2 y = 3x – 1 2.) –4x + 3y = –12 –2x + 3y = –18 Will someone help me and explain this to me?

OpenStudy (anonymous):

@x1xofxaxkndx can u help me pleaseeee

OpenStudy (anonymous):

you want the solution??

OpenStudy (anonymous):

1.) y = –x + 2 y = 3x – 1 2.) –4x + 3y = –12 –2x + 3y = –18

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

i so wish i could help but im dumb when it comes to this

OpenStudy (anonymous):

3x + 2y = 10 ----> equation(1) 2x - 3y = 12 ----> equation(2) multiply equation(1) by -2 and equation (2) by 3 -6x - 4y = -20 6x - 9y = 36 -------------------- -13y = 16 y = -16/13 Substitute y = -16/13 into equation(1) or equation(2) to solve for x, 3x + 2y = 10 3x + 2(-16/13) = 10 3x - 32/13 = 10 39x - 32 = 130 39x = 162 x= 162/39 = 54/13

OpenStudy (anonymous):

thats how to get the solution

OpenStudy (anonymous):

okay so how would i graph those?

OpenStudy (anonymous):

thanks anyways sweetie (: @x1xofxaxkndx

OpenStudy (anonymous):

im not understanding?

OpenStudy (anonymous):

@growdabone could u help me?

OpenStudy (anonymous):

i dont know how to graph

OpenStudy (anonymous):

what are you wanting explained?

OpenStudy (anonymous):

how to solve both of them and graph them!

OpenStudy (anonymous):

www.wolframalpha.com type in both your equations and type plot you will get the graphs

OpenStudy (anonymous):

i dont get graphing that well... but if u still want i can try to work with you on the problem, also freemathhelp.com will graph it for u, u gotta go to algebra and then the algebra calculator!

OpenStudy (anonymous):

Thank you every one (: